Shift report 20011228 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Dec 28 2001 - 16:06:34 EST

  • Next message: Flemming Videbaek: "The lasst weekend in 2001 (fwd)"

    ------------------------------------------------------------
    Date: Fri Dec 28 16:06:33 2001, Shift: 08:00-16:00
    Supervisor: D. Beavis, with: 
    Log book#11 pages 104 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    At beginning of shift there was the remains of a store.
    Stopped taking data at 8:30
    Did a ped run and a dvm run.
    Next store started at 1200 and lasted till 1530 with
    unpolarized beam.
    Run 6186 had 122k trig 1, 5900 trig 3, 159k trig 5 and 20k trig 6
    A new stored with polarized protons is being prepared.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Dec 28 2001 - 16:07:21 EST